# AC-Lite : A Lightweight Image Captioning Model for Low-Resource Assamese Language

### v1

@@ Abstract
Neural networks have significantly advanced AI applications, yet their real-world adoption remains constrained by high computational demands, hardware limitations, and accessibility challenges. In image captioning, many state-of-the-art models have achieved impressive performances while relying on resource-intensive architectures. This made them impractical for deployment on resource-constrained devices. This limitation is particularly noticeable for applications involving low-resource languages. We demonstrate the case of image captioning in Assamese language, where lack of effective, scalable systems can restrict the accessibility of AI-based solutions for native Assamese speakers. This work presents AC-Lite, a computationally efficient model for image captioning in low-resource Assamese language. AC-Lite reduces computational requirements by replacing computation-heavy visual feature extractors like FasterRCNN with lightweight ShuffleNetv2x1.5. Additionally, Gated Recurrent Units (GRUs) are used as the caption decoder to further reduce computational demands and model parameters. Furthermore, the integration of bilinear attention enhances the model’s overall performance. AC-Lite can operate on edge devices, thereby eliminating the need for computation on remote servers. The proposed AC-Lite model achieves 82.3 CIDEr score on the COCO-AC dataset with 1.098 GFLOPs and 25.65M parameters.
@@ 1 Introduction
Deep neural networks have elevated the capability of machines for image understanding and natural language modeling. Automatic Image Captioning (AIC) is a research area that enables machines to automatically generate meaningful textual descriptions for input images. This capability is essential for multimodal AI applications, including assistive technologies for visually impaired individuals, intelligent human-machine interfaces, and visual content retrieval [1]. However, despite significant advancements in image captioning research  [2, 3, 4, 5, 6], the practical deployment of the state-of-art models remain a major challenge. This can be attributed to the prioritization of achieving arbitrarily high overall accuracy at the cost of computational efficiency. This often renders most models unsuitable for certain real-world applications that require on-device processing and low-latency inference on energy-efficient edge devices.
Table 1: Component-wise computational complexity (in Million FLOPs) and parameter (in Million) comparison of various deep learning-based Image Captioning models. Note that the FLOPs and Parameter calculations are based on the hyper-parameters mentioned in the respective papers.
Model Catagories Visual Feature Encoder Visual Feature Refiner Decoder With Attention Total Model MFLOPs Param (M) Model MFLOPs Param (M) Model MFLOPs Param (M) MFLOPs Param (M) Show and Tell [3] GoogleNet 1500 6.6 - - - LSTM 702.207 19.506 2202.207 26.106 Show attend and tell [2] Vgg16 15470 138.4 - - - LSTM 722.004 20.056 16192.004 158.456 BUTD [4] FasterRCNN 117330 63.63 - - - bottom-up attention + LSTM 2122.919 58.97 119452.919 122.6 AoANet [5] FasterRCNN 117330 63.63 AoA 285.554 37.79 AoA + LSTM 1614.685 17.832 119230.239 119.252 ORT [6] FasterRCNN 117330 63.63 ORT Encoder 955.404 19.923 Transformer Decoder 17511.8 35.794 135797.204 119.347
Early image captioning models have used template-based and retrieval-based approaches to generate captions. These early approaches have either filled blank spaces in a fixed sentence template or retrieved similar captions from a prefix caption database. Though these approaches were simple, the generated captions were of fixed size, or the captions were often irrelevant to image content. The introduction of deep learning-based image captioning models could mitigate these issues up to some extent [7]. The deep learning-based models mainly consist of three main components – (a) Visual Feature Encoder, (b) Visual Feature Refiner, and (c) Decoder With Attention (see Figure 1). These models [2, 3, 4, 5, 6] have achieved impressive performance by leveraging computation-heavy visual feature extractors. However, these advancements came with a significant increase in computational demands and memory footprint. Table 1 highlights the progression of notable deep learning-based image captioning models along with computations (measured in MFLOPs or Million FLOPs) and parameter counts (in Million). As shown in Table 1 models like Show and Tell [3] and Show Attend and Tell [2] used simpler architectures and achieved comparable performance. However, these models used a heavy Convolutional Neural Network (CNN) such as VGG16 [36] and GoogleNet [8] as a visual feature encoder, which increased the overall computation. Additionally, models like BUTD [4] and AoANet [5] relied on FasterRCNN [9] as the visual feature encoder. The computation-heavy FasterRCNN contributed significantly to the overall computation.
Due to the transformer-based architectures, the computational requirements for the ORT [6] model surpassed others. The high computational demand and large parameter counts made it infeasible for real-time caption generation on resource-constrained devices such as smartphones, tablets, and embedded AI systems. These high computational costs not only hinder on-device deployment but also increase reliance on cloud-based inference, leading to latency issues and higher energy consumption. Such dependencies create accessibility barriers, particularly in remote areas where internet connectivity is unreliable. This observation motivates the present work to develop a lightweight image captioning model by utilizing lightweight CNNs as visual feature encoders.
@@
Most image captioning research has been centered around high-resource languages such as English, Chinese [10], and German [11], while low-resource languages like Assamese have received minimal attention. The Assamese language is the official language of the North-East Indian state of Assam. It is an Indo-Aryan language similar to Hindi, Bengali, and Oriya. The Assamese language follows a left-to-right writing script similar to the Bengali language. Assamese is a morphologically rich language with complex grammar having inflection, gender and tense markers. Additionally, Assamese has 15.31 million native speakers, out of which  85% reside in rural areas and are more fluent in Assamese than Englishwww.language.census.gov.in/showLanguageCensusData. A lightweight Assamese image captioning system for edge devices could address real-world challenges. An AI-based assistive technology using Assamese image captioning can help 0.3% of the population of Assam who suffer from visual impairment www.finance.assam.gov.in. Moreover,  70% of students from Assam get their elementary education in the Assamese language www.elementary.assam.gov.in/. Therefore, a computationally efficient Assamese captioning model can help in smart education technology for language learning [12]. The design of image captioning models with lower computations and parameters is vital for such practical applications. This is particularly applicable for other low-resource languages as well for broader accessibility and usability of AI systems.
@@
While deep learning-based image captioning models have achieved remarkable success, their adoption in real-world applications remains constrained by high computational demands and hardware limitations. Most state-of-the-art models prioritize performance, often at the expense of efficiency, making them impractical for deployment on edge devices. This limitation is particularly evident in applications like assistive technologies for visually impaired individuals, where computational efficiency is crucial for real-time operation. To address these challenges, this work introduces AC-Lite, a lightweight image captioning model for low-resource Assamese language. AC-Lite is designed for deployment on computation resource-constrained devices (e.g. smartphones). Additionally, the model should operate smoothly on the device hardware without connecting to any remote server for caption generation. To achieve this, the proposed AC-Lite model uses computationally efficient CNN models for the visual feature encoder. Additionally, the Gated Recurrent Units (GRU) have been employed in the decoder due to their fewer parameters, fewer computational requirements, and faster training times. Further, bilinear attention [13] is also employed with the proposed AC-Lite to boost the performance. Moreover, a reinforcement learning-based approach [14] is applied for further improvement of captioning performance. The performance of AC-Lite is evaluated by using COCO-Assamese Caption (COCO-AC) and the Flickr30K-Assamese Caption (Flickr30K-AC) datasets proposed in [13]. The major contributions of this work are as follows.
@@ 2 Related Work
Transformer architectures have improved image captioning performance by visual feature refinement and by enhancing the semantic accuracy of captions. Cornia et al. [15] proposed a meshed-memory transformer, linking encoder and decoder layers for multi-level vision-language relationships, with a memory-augmented attention mechanism. Pan et al. [16] introduced X-linear attention in transformers, enabling second-order interactions for richer representations. Herdade et al. [6] adjusted attention weights using geometric shapes and spatial positions. Nair et al. [17] replaced trigonometric positional embeddings with spatial positions of salient regions. However, these transformer-based architectures are quite computation intensive and cannot be deployed on edge devices (see Table 1).
@@
Lightweight Convolutional Neural Networks have gained popularity for their operational efficiency on computation resource-constrained devices. For instance, MobileNetV1 [18] and MobileNetV2 [19] used techniques like depth-wise separable convolutions and inverted bottleneck blocks for reducing computations. Following that, other lightweight CNNs like ShuffleNetV1 [29], ShuffleNetV2 [21], EfficientNet [22], and MobileNetV3 [20] emerged. These lightweight CNNs also have scaled-down versions like ShuffleNetV2X0.5, ShuffleNetV2X1.0, ShuffleNetV2X1.5, ShuffleNetV2X2.0 and MobileNetv3_small which operate in different configurations. However, these CNNs favor speed over accuracy, leading to slightly reduced performance.
@@
Lightweight Image Captioning models focus on efficiency for resource-constrained settings. Models like ACORT [25] and COMIC [30] used parameter and vocabulary size reduction techniques. ACORT [25] reduced parameters through cross-layer and attention parameter sharing, achieving a 13$\times$ reduction compared to transformers [31]. Additionally, the authors introduced Radix Encoding to minimize vocabulary size. The Radix Encoding method encodes the original vocabulary into a smaller set of symbols, thus reducing the parameter count in the embedding matrix. COMIC [30] employed a similar vocabulary encoding technique to reduce parameters. LCM-Captioner [23] used TextLighT to map visual features to a lower-dimensional space. Zhang et al. substituted FasterRCNN with MobileNetV3 for visual encoding and employed the transformer decoder for caption decoding. LightCap [24] used pre-trained CLIP model for visual feature encoding without relying on resource-intensive object detectors. SMALLCAP [26] combined retrieval-augmented generation with pre-trained CLIP vision encoder and GPT-2, thereby reducing total trainable parameters. I-Tuning [27] used cross-attention to link the frozen CLIP and GPT-2 models, thus reducing trainable parameters by up to 10$\times$ while maintaining the overall performance.
@@
Image Captioning in Indian Languages like Assamese, Bengali, and Hindi is scarce, primarily due to the lack of large-scale image-caption datasets. Mishra et al. [33] introduced MSCOCO-Hindi by translating MSCOCO-English captions into Hindi. Additionally, the authors employ ResNet101 [35] as visual feature encoder and GRU with attention for caption generation. In another work, Mishra et al. [32] used the transformer decoder with ResNet101 for Hindi caption generation. In Assamese language, Nath et al. [34] used a similar translation-based approach to create Assamese image-caption dataset and employed EfficientNetB3 with GRU for caption generation. However, automatically translated Assamese captions often suffer from syntactic and semantic errors [13] and that is not addressed in [34]. Choudhury et al. [13] addressed this issue by manually correcting the translation errors to develop the COCO-Assamese and Flickr30K-Assamese datasets. Additionally, the authors employed FasterRCNN for visual feature encoding and Bi-LSTM with bilinear attention for caption generation. Further, Choudhury et al. [28] showed that models trained exclusively on Assamese image-caption datasets outperform those trained on English datasets and subsequently translated into Assamese. This highlights the importance of developing image caption datasets and models for low-resource languages like Assamese.
@@ 3 Proposed Work
The proposed AC-Lite model consists of three main components – (a) Visual Feature Encoder, (b) Bilinear Attention, and (c) Gated Recurrent Unit (GRU) for caption decoding as shown in Figure 2(a). AC-Lite uses pre-trained CNNs (trained on ImageNet) as the visual feature encoder. Initially, the Visual Feature Encoder takes an input image and transforms it into a feature map of size $d_{a}\times n_{h}\times n_{w}$. These feature maps are then flattened to create visual feature vectors $\mathbf{A}=[\mathrm{\boldsymbol{a}}_{1},\ldots\mathrm{\boldsymbol{a}}_{i},\ldots\mathrm{\boldsymbol{a}}_{n_{a}}]$ ($\mathrm{\boldsymbol{a_{i}}}\in\mathbb{R}^{d_{a}\times 1},\mathbf{A}\in\mathbb{R}^{d_{a}\times n_{a}},n_{a}=n_{h}\times n_{w}$). The extracted visual features $\mathbf{A}$ are then fed to the Bilinear Attention module to generate the attended visual feature $\boldsymbol{\hat{a}}_{t}$ at each decoding step $t$.
The Bilinear Attention module has shown superior performance for Assamese image captioning task [13]. In the original paper [13], the bilinear attention module consists of an attention LSTM $\mathbf{LSTM_{A}}$ and attention weight computation mechanism. However, to reduce computation, we replace $\mathbf{LSTM_{A}}$ with a Gated Linear Unit (GRU) $\mathbf{GRU_{A}}$ as shown in Figure 2(b). At the $t^{th}$ decoding step, $\mathbf{GRU_{A}}$ produces a partial caption, and its context is captured in $\mathbf{h}_{t}^{A}\in\mathbb{R}^{d_{h}\times 1}$. $\mathbf{GRU_{A}}$ takes mean pooled visual features $\bar{{\boldsymbol{a}}}=\frac{\sum_{i=1}^{n_{a}}\boldsymbol{a}_{i}}{n_{a}}$ concatenated with previous hidden state $\mathbf{h}_{t-1}^{G}$ of the decoder GRU as input $\mathbf{x}_{t-1}=\left[\mathbf{h}_{t-1}^{G}:\bar{{\boldsymbol{a}}}\right]$ to produce the partial caption.

### v2

@@ Abstract
Most existing works in image caption synthesis use computation heavy deep neural networks and generates image descriptions in English language. This often restricts this important assistive tool for widespread use across language and accessibility barriers. This work presents AC-Lite, a computationally efficient model for image captioning in low-resource Assamese language. AC-Lite reduces computational requirements by replacing computation-heavy deep network components with lightweight alternatives. The AC-Lite model is designed through extensive ablation experiments with different image feature extractor networks and language decoders. A combination of ShuffleNetv2x1.5 with GRU based language decoder along with bilinear attention is found to provide the best performance with minimum compute. AC-Lite was observed to achieve an 82.3 CIDEr score on the COCO-AC dataset with 2.45 GFLOPs and 22.87M parameters.
@@ 1 Introduction
Deep neural networks have elevated the capability of machines for image understanding and natural language modeling. Automatic Image Captioning (AIC) is a research area that enables machines to automatically generate meaningful textual descriptions for input images. This capability is essential for multimodal AI applications, including assistive technologies for visually impaired individuals, intelligent human-machine interfaces, and visual content retrieval [1]. However, despite significant advancements in image captioning research  [2, 3, 4, 9, 10], the practical deployment of the state-of-art models remain a major challenge. This can be attributed to the prioritization of achieving arbitrarily high overall performance at the cost of computational efficiency. This often renders most models unsuitable for certain real-world applications that require on-device processing and low-latency inference on energy-efficient edge devices.
@@
Early image captioning models have used template-based and retrieval-based approaches to generate captions. These early approaches have either filled blank spaces in a fixed sentence template or retrieved similar captions from a prefix caption database. Though these approaches were simple, the generated captions were of fixed size, or the captions were often irrelevant to the image content. The introduction of deep learning-based image captioning models could mitigate these issues up to some extent [11]. The deep learning-based models mainly consist of three main components – (a) Visual Feature Encoder, (b) Visual Feature Refiner, and (c) Decoder With Attention (see Figure 1) [11]. The deep learning based models [2, 3, 4, 9, 10] have achieved impressive performance by leveraging computation-heavy visual feature extractors. However, these advancements came with a significant increase in computational demands and memory footprint. Table 1 highlights notable deep learning-based image captioning models along with computations (measured in Billion FLOPs or GFLOPs) and parameter counts (in Millions)FLOPs and Parameters are computed using fvcore library (https://github.com/facebookresearch/fvcore).. As shown in Table 1, models like Show and Tell [3] and Show Attend and Tell [2] used simpler architectures and achieved comparable performance. However, models like Show Attend and Tell used a heavy Convolutional Neural Network (CNN) such as VGG19 [40] as the visual feature encoder, which increased the overall computation. Additionally, models like BUTD [4] and AoANet [9] relied on FasterRCNN [13] as the visual feature encoder. The computation-heavy FasterRCNN contributed significantly to the overall computation.
Table 1: Component-wise computational complexity (in Giga FLOPs) and parameter (in Millions) comparison of various deep learning-based Image Captioning models. Note that the FLOPs and Parameter calculations are based on the hyper-parameters mentioned in the respective papers.
Model Categories Visual Feature Extractor Visual Feature Refiner Decoder with Attention Total Model GFLOPs Params (M) Model GFLOPs Params (M) Model GFLOPs Params (M) GFLOPs Params (M) Show and Tell GoogLeNet 1.5 6.6 - - - LSTM 0.131 13.4 1.631 20 Show, Attend and Tell VGG19 19.5 20 - - - LSTM 0.401 14.45 19.901 34.45 BUTD FasterRCNN 117.33 63.63 - - - UpDown 0.945 52.12 118.275 115.75 AoANet FasterRCNN 117.33 63.63 AoA refiner 4.527 44.09 AoA Decoder 0.586 41.48 122.443 149.2 ORT FasterRCNN 117.33 63.63 ORT Encoder 1.98 18.92 ORT Decoder 0.795 54.92 120.105 137.47 M2Transformer FasterRCNN 117.33 63.63 Memory Augmented Encoder 1.093 10.6 Meshed Decoder 0.974 38.44 119.397 112.67 Transformer FasterRCNN 117.33 63.63 Transformer Encoder 1.949 18.92 Transformer Decoder 0.916 36 120.195 118.55
Due to the transformer-based architectures, the computational requirements for models like ORT [10] and M2 [19] are usually higher than other models. Real time transformer model-based image caption synthesis on resource constrained embedded devices is often challenged by high computations and large number of parameters. These high computational costs not only hinder on-device deployment but also increase reliance on cloud-based inference. This leads to latency issues and higher energy consumption. Moreover, such dependencies create accessibility barriers, particularly in remote areas where internet connectivity is unreliable. This observation motivates the present work to develop a lightweight image captioning model by utilizing low latency CNNs as visual feature encoders.
@@
Most image captioning research has been centered around high-resource languages such as English, Chinese [14], and German [15], while low-resource languages like Assamese have received minimal attention. The Assamese language is the official language of the North-East Indian state of Assam. It is an Indo-Aryan language similar to Hindi, Bengali, and Oriya. The Assamese language follows a left-to-right writing script similar to the Bengali language. Assamese is a morphologically rich language with complex grammar having inflection, gender and tense markers. Additionally, Assamese has 15.31 million native speakers, out of which  85% reside in rural areas and are more fluent in Assamese than Englishwww.language.census.gov.in/showLanguageCensusData. A lightweight Assamese image captioning system for edge devices could address real-world challenges like assisting the visually impaired. An AI-based assistive technology using Assamese image captioning can help 0.3% of the population of Assam who suffer from visual impairment www.finance.assam.gov.in. When deployed on smartphones, such a system enables on-device accessibility without relying on constant internet connectivity, making it a practical and inclusive solution for visually impaired users in resource-constrained settings. Moreover,  70% of students from Assam get their elementary education in the Assamese language www.elementary.assam.gov.in/. Therefore, a computationally efficient Assamese captioning model can help in smart education technology for language learning [16]. The design of image captioning models with lower computations and parameters is vital for such practical applications. This is particularly applicable for other low-resource languages as well for broader accessibility and usability of AI systems.
@@
While deep learning-based image captioning models have achieved remarkable success, their adoption in real-world applications remains constrained by high computational demands and hardware limitations. Most state-of-the-art models prioritize performance, often at the expense of efficiency. This makes them impractical for deployment on edge devices. This limitation is particularly evident in applications like assistive technologies for visually impaired individuals, where computational efficiency is crucial for real-time operation. To address these challenges, this work introduces AC-Lite, a lightweight image captioning model for low-resource Assamese language. Unlike existing approaches, this is the first effort specifically aimed at developing a computationally efficient captioning model for Assamese. AC-Lite is designed for deployment on computation resource-constrained devices (e.g. smartphones). Additionally, the model should operate smoothly on the device hardware without connecting to any remote server for caption generation. To achieve this, the proposed AC-Lite model uses computationally efficient CNN models for the visual feature encoder. Moreover, the Gated Recurrent Units (GRU) have been employed in the decoder due to their fewer parameters, fewer computational requirements, and faster training times. Further, bilinear attention [17] is also employed with the proposed AC-Lite to boost the performance. Finally, a reinforcement learning-based approach [18] is applied for further improvement of captioning performance. The performance of AC-Lite is evaluated by using COCO-Assamese Caption (COCO-AC) and the Flickr30K-Assamese Caption (Flickr30K-AC) datasets proposed in [17]. The major contributions of this work are as follows.
@@ 2 Related Work
Transformer architectures have improved image captioning performance by visual feature refinement and by enhancing the semantic accuracy of captions. Cornia et al. [19] proposed a meshed-memory transformer, linking encoder and decoder layers for multi-level vision-language relationships, with a memory-augmented attention mechanism. Pan et al. [20] introduced X-linear attention in transformers, enabling second-order interactions for richer representations. Herdade et al. [10] adjusted attention weights using geometric shapes and spatial positions. Nair et al. [21] replaced trigonometric positional embeddings with spatial positions of salient regions. Notably, these transformer-based architectures are quite computation intensive (see Table 1).
@@
Lightweight Convolutional Neural Networks have gained popularity for their operational efficiency on resource-constrained devices. For instance, MobileNetV1  [22] and MobileNetV2 [23] used techniques like depth-wise separable convolutions and inverted bottleneck blocks for reducing computations. Following that, other lightweight CNN architectures like ShuffleNetV1 [33], ShuffleNetV2 [25], EfficientNet [26], and MobileNetV3 [24] were proposed. These lightweight CNNs also have their scaled-down versions like ShuffleNetV2X0.5, ShuffleNetV2X1.0, ShuffleNetV2X1.5, ShuffleNetV2X2.0 and MobileNetv3_small which operate in different configurations. However, a trade-of between performance and computations is observed across all architectures.
@@
Lightweight Image Captioning models focus on efficiency for resource constrained settings. Models like ACORT [29] and COMIC [34] used parameter and vocabulary size reduction techniques. ACORT [29] reduced parameters through cross-layer and attention parameter sharing, achieving a 13$\times$ reduction compared to transformers [35]. Additionally, the authors introduced Radix Encoding to minimize vocabulary size. The Radix Encoding method encodes the original vocabulary into a smaller set of symbols, thereby reducing the parameter count in the embedding matrix. COMIC [34] employed a similar vocabulary encoding technique to reduce parameters. LCM-Captioner [27] used TextLighT to map visual features to a lower-dimensional space. Zhang et al. [8] substituted FasterRCNN with MobileNetV3 for visual encoding and employed the transformer decoder for caption decoding. LightCap [28] used pre-trained CLIP model for visual feature encoding without relying on resource-intensive object detectors. SMALLCAP [30] combined retrieval-augmented generation with pre-trained CLIP vision encoder and GPT-2, thereby reducing total trainable parameters. I-Tuning [31] used cross-attention to link the frozen CLIP and GPT-2 models, thus reducing trainable parameters by up to 10$\times$ while maintaining the overall performance. However, these lightweight image captioning models rely on English pre-trained language models, making them unsuitable for low-resource languages like Assamese. This motivates the need to develop lightweight captioning models for low-resource languages like Assamese.
@@
Image Captioning in Indian Languages like Assamese, Bengali, and Hindi is scarce, primarily due to the lack of large-scale image-caption datasets. Mishra et al. [37] introduced MSCOCO-Hindi by translating MSCOCO-English captions into Hindi. Additionally, the authors employ ResNet101 [39] as visual feature encoder and GRU with attention for caption generation. In another work, Mishra et al. [36] used the transformer decoder with ResNet101 for Hindi caption generation. In Assamese language, Nath et al. [38] used a similar translation-based approach to create Assamese image-caption dataset and employed EfficientNetB3 with GRU for caption generation. However, automatically translated Assamese captions often suffer from syntactic and semantic errors [17] and that is not addressed in [38]. Choudhury et al. [17] addressed this issue by manually correcting the translation errors to develop the COCO-Assamese and Flickr30K-Assamese datasets. Additionally, the authors employed FasterRCNN for visual feature encoding and Bi-LSTM with bilinear attention for caption generation. In a related study, Choudhury et al. [5] incorporated semantic attributes with visual features to improve Assamese captioning performance. The semantic attributes are key concepts presented in an image, such as object and place names, attributes, and actions. These semantic attributes are mapped to feature vectors using FastText [7] embeddings trained on a large Assamese corpus. FastText was chosen for its subword-level modeling, which captures the inflectional properties of Assamese effectively. Further, Choudhury et al. [32] demonstrated that models trained directly on Assamese datasets performed better than those trained on English datasets followed by translation into Assamese. This highlights the importance of developing image caption datasets and models for low-resource languages like Assamese.
@@ 3 Proposed Work
The proposed AC-Lite model consists of three main components – (a) Visual Feature Encoder, (b) Bilinear Attention, and (c) Gated Recurrent Unit (GRU) for caption decoding as shown in Figure 2. AC-Lite uses pre-trained CNNs (trained on ImageNet) as the visual feature encoder. Initially, the Visual Feature Encoder takes an input image and transforms it into a feature map of size $d_{a}\times n_{h}\times n_{w}$. These feature maps are then flattened to create visual feature vectors $\mathbf{A}=[\mathrm{\boldsymbol{a}}_{1},\ldots\mathrm{\boldsymbol{a}}_{i},\ldots\mathrm{\boldsymbol{a}}_{n_{a}}]$ ($\mathrm{\boldsymbol{a_{i}}}\in\mathbb{R}^{d_{a}\times 1},\mathbf{A}\in\mathbb{R}^{d_{a}\times n_{a}},n_{a}=n_{h}\times n_{w}$). The extracted visual features $\mathbf{A}$ are then fed to the Bilinear Attention module to generate the attended visual feature $\boldsymbol{\hat{a}}_{t}$ at each decoding step $t$.
The Bilinear Attention module has shown superior performance for Assamese image captioning task [17]. In the original paper [17], the bilinear attention module consists of an attention LSTM $\mathbf{LSTM_{A}}$ and attention weight computation mechanism. However, to reduce computation, the $\mathbf{LSTM_{A}}$ in the bilinear attention module is replaced with a Gated Linear Unit (GRU) $\mathbf{GRU_{A}}$ as shown in Figure 2. At the $t^{th}$ decoding step, $\mathbf{GRU_{A}}$ produces a partial caption, and its context is captured in $\mathbf{h}_{t}^{A}\in\mathbb{R}^{d_{h}\times 1}$. $\mathbf{GRU_{A}}$ takes mean pooled visual features $\bar{{\boldsymbol{a}}}=\frac{\sum_{i=1}^{n_{a}}\boldsymbol{a}_{i}}{n_{a}}$ concatenated with previous hidden state $\mathbf{h}_{t-1}^{G}$ of the decoder GRU as input $\mathbf{x}_{t-1}=\left[\mathbf{h}_{t-1}^{G}:\bar{{\boldsymbol{a}}}\right]$ to produce the partial caption.